OpenStudy (anonymous):

Determine whether the graphs of the two equations are perpendicular. x+2y=7 2x+4y=5 Yes or no?

OpenStudy (zarkon):

no

myininaya (myininaya):

so it will be easy to tell if we write both in y=mx+b form so first line we have x+2y=7 2y=-x+7 y=-x/2+7/2 second one we have 2x+4y=5 4y=-2x+5 y=-2x/4+5/4 y=-x/2+5/4 so the lines are parallel